# To appear on arXiv arXiv:2507.12345 [cs.LG] Version 1, July 2025

### v1

@@ 1. Introduction and Preliminaries
A pairwise affinity matrix $A$ is a square matrix (size $N\times N$ for $N$ elements) where each entry $A_{ij}$ encodes a relationship or similarity between element $i$ and element $j$. Such matrices have long been used in machine learning and pattern recognition to represent graphs or networks of interactions. Early examples include graph adjacency matrices in spectral clustering (where $A_{ij}$ might be a Gaussian similarity between data points) and social network graphs for centrality analysis. The key idea is that by comparing all pairs of elements, one can capture global structure or importance. This concept underpins modern self-attention mechanisms: self-attention computes a learned affinity matrix between components of the input (e.g., between tokens in a sequence) and uses it to decide how information flows between those components.
Self-attention (also called intra-attention) refers to an attention mechanism that relates a sequence or set with itself, as opposed to the earlier “attention” in sequence-to-sequence models which related one sequence with another (e.g., decoder attending to encoder states in machine translation [6]). Importantly, self-attention is implemented via an affinity matrix $A$ of size $N\times N$ (for $N$ input elements) that effectively forms a fully-connected graph over the elements. Each $A_{ij}$ indicates how much element $i$ should pay attention to element $j$. This structural idea – computing pairwise comparisons to inform a weighting – has a rich history across domains, even if it wasn’t always called “attention.”
@@
Below is a chronological overview of influential works across domains that utilized the concept of pairwise affinity matrices, culminating in the modern self-attention framework:
  • 1998–2005 (Early Vision & Graph Ideas): Bilateral Filter (1998) and Non-Local Means (2005) in image processing — used fixed Gaussian affinity between pixels for smoothing [7, 8]. PageRank (1998) in graph theory — used the web link graph (adjacency matrix) to rank webpages, analogous to ranking nodes by an infinite walk (conceptually similar to Inf-FS’s approach on a feature graph) [9]. While these methods were not neural or learnable, they highlighted the power of pairwise affinity matrices and propagation over graphs (e.g., PageRank as a stationary distribution of an infinite walk; bilateral filtering as a one-step weighted average).
  • 2014 (Neural Feature Selection): Qian Wang et al., “Attentional Neural Network” (NeurIPS 2014) — one of the early works to incorporate a learned attention mechanism for feature selection in neural networks [19]. Top-down attention signals modulate bottom-up feature activations, learning to emphasize relevant inputs.
  • 2015 (Graph-Based Feature Selection): Giorgio Roffo et al., “Infinite Feature Selection (Inf-FS)” (ICCV 2015) — proposed a framework for selecting features based on a fully connected graph of affinities [1]. The method computes a ranking score via an infinite matrix power series:
@@
    where $A$ is an affinity matrix capturing feature–feature relationships. Importantly, $A$ is not fixed: the authors explicitly state it can be handcrafted or learned depending on the application. This positions Inf-FS as a general framework for reasoning over pairwise affinities, which can subsume self-attention as a specific case where $A$ is computed via learned dot-product similarity. In this view, the distinction between Inf-FS and self-attention lies in the parameterization and operational use of $A$, not in the underlying structure.
  • 2015 (Neural Machine Translation Attention): Bahdanau, Cho, Bengio, “Neural Machine Translation by Jointly Learning to Align and Translate” (ICLR 2015) — introduced neural attention for aligning source and target sequences [6], with soft alignment weights:
    $e_{ij}=a(s_{i-1},h_{j}),\quad\alpha_{ij}=\frac{\exp(e_{ij})}{\sum_{k}\exp(e_{ik})}$
    paving the way for broader attention mechanisms.
  • 2016 (Self-Attention in RNNs): Cheng, Dong, Lapata (ACL 2016) [20] implemented self-attention within RNNs. Parikh et al. (2016, EMNLP) [22], Lin et al. (ICLR 2017) [23], and Paulus et al. (2017) [21] further explored attention for sentence embeddings and summarization, showing that attention can augment or replace recurrent computation.
  • 2017 (Attention for Sets and Graphs): Santoro et al., “Relation Networks” (NeurIPS 2017) — introduced relational reasoning via pairwise interactions across sets [24]. Lee et al., “Set Transformers” (2019) [25] extended this using self-attention to model unordered sets, leveraging the permutation-invariance of affinity matrices.
  • 2017 (Inf-FS with Learning): Roffo et al., “Infinite Latent Feature Selection” (ICCV 2017) — extended Inf-FS by learning the affinity matrix from data using probabilistic models [2]. This learning capability further bridges Inf-FS with neural attention. Related works include Roy et al. (2018) [26], Gui et al. (2019) [27], and Abid et al. (2019) [28], which use attention-like mechanisms for feature selection.
  • 2017 (Transformer and Self-Attention): Vaswani et al., “Attention Is All You Need” (NeurIPS 2017) — introduced the Transformer architecture and scaled dot-product attention:
@@
    [5]. This represented a major shift, replacing recurrence with self-attention as the central mechanism for sequence modeling.
  • 2018 (Non-Local Neural Networks): Wang et al., “Non-Local Neural Networks” (CVPR 2018) — extended self-attention to vision by computing long-range dependencies in space and time via pairwise feature similarities [30].
  • 2018 (Graph Attention Networks): Veličković et al., “Graph Attention Networks” (ICLR 2018) — applied attention to graph-structured data, learning task-specific affinities over graph edges [31].
  • 2019–2020 (Efficiency): To scale attention to longer sequences, efficient approximations were developed: Reformer [32], Linformer [33], Performer [34], and Efficient Attention [35]. These methods reduce memory or compute costs while preserving the core pairwise interaction model.
  • 2021 (Inf-FS in TPAMI): Roffo et al. published the journal version of Inf-FS in TPAMI [3], explicitly positioning the method in relation to attention by referencing concepts like feature reweighting. Ramsauer et al. (ICLR 2021) [37] also drew formal connections between Hopfield networks and Transformers, interpreting attention as memory-based retrieval over affinity matrices.
  • 2022 and Beyond: Transformers are now pervasive across domains: AlphaFold (protein folding), CLIP (vision-language), Switch Transformers (scalable sparse attention). Affinity-based mechanisms are also used in clustering, segmentation, and structured modeling, with ongoing research into sparsity, axial attention, and learned kernel structures.
@@ 2. Graph-Based Feature Selection and Early Affinity Approaches
Fast-forward to 2018: Xiaolong Wang et al. published Non-Local Neural Networks [30], explicitly inspired by the self-attention idea in NLP and by classical non-local filtering in vision. They proposed a non-local block as a generic module that can be inserted into convolutional neural networks to allow global dependencies to be captured. In a non-local block, the feature map (say of shape $N$ locations by $C$ channels) is treated similarly to a sequence: they compute an affinity between any two locations $i$ and $j$ in the feature map and use it to adjust the features. In the simplest version, they define $f(x_{i},x_{j})=\theta(x_{i})^{T}\phi(x_{j})$, where $x_{i}$ is the feature vector at position $i$ (like a pixel or a region), and $\theta,\phi$ are learnable linear projections (analogous to queries and keys) [30]. This is exactly the dot-product similarity from self-attention (sometimes they also experimented with Gaussian $f(x_{i},x_{j})=e^{x_{i}^{T}x_{j}}$ or a concatenation-based function) [30]. They then apply a softmax normalization over $j$ for each $i$ (for the Gaussian version; for the pure dot-product version they normalized by $1/N$) [30]. Finally, they compute an output at $i$ as $y_{i}=\sum_{j}f(x_{i},x_{j}),g(x_{j})$, where $g(x_{j})$ is typically a linear embedding (like the “value” in attention, noted as $W_{g}x_{j}$) [30]. They add $y$ back to the input (residual connection) and feed it into the next layer. This non-local operation can be applied to image features (where $i,j$ index spatial positions) or video (indexing spatial-temporal positions) or even more abstract graph-like data.
@@
3. Extension of Inf-FS Across Applications
The affinity matrix $A$ introduced in Infinite Feature Selection (Inf-FS) has played a central role across a range of applications and follow-up work. Many subsequent papers have extended, adapted, or leveraged the $A$ matrix structure for different learning contexts, validating its generality and versatility. One notable extension was the use of eigenvector centrality as a scoring mechanism in feature graphs. In [10], features are ranked using the principal eigenvector of the affinity matrix, highlighting the most influential nodes in the graph. This direction aligns with the broader understanding of graph centrality and its role in importance estimation.
The importance of affinity-based feature relevance is also emphasized in [11], which examines how ranking and learning interact in various pattern recognition tasks. That work conceptualizes ranking not just as an output but as an operational tool within learning pipelines, further motivating the use of matrices like $A$ for ordering features. In practical applications, such as visual tracking, Inf-FS and its variants were adapted to work in online scenarios, dynamically selecting features over time [12]. In a biomedical context, Inf-FS was used on neuroimaging data to identify biomarkers indicative of stroke recovery [13], showing that the method generalizes to structured medical data and can model complex biological interactions. Recent works have also explored the intersection of feature selection with large language models and attention gates. The concept of self-attention has been further connected to feature selection in models like those in [14], [15], and [16], which incorporate gradient routing and feature-level gating to perform instance-specific, end-to-end selection. These developments further confirm the convergence of affinity matrix modeling and attention-based architectures.
For users and researchers, Infinite Feature Selection (Inf-FS) has also been implemented in public libraries. In addition to the MATLAB-based Feature Selection Library (FSLib) [17], [18], a Python version is available as PyIFS on PyPI (https://pypi.org/project/PyIFS/), enabling broader experimentation and integration into modern deep learning pipelines.
@@ Conclusion
Acknowledgements
We thank the authors of the original Inf-FS paper and the broader community working on self-attention, feature selection, and graph-based learning, whose ideas we explore and unify in this work.

### v2

# To appear on arXiv arXiv:2507.14560 [cs.LG] Version 1, July 2025

@@ 1. Introduction and Preliminaries
A pairwise affinity matrix $A\in\mathbb{R}^{N\times N}$ represents relationships or similarities between elements $i,j\in\{1,\dots,N\}$. Such matrices are fundamental in pattern recognition, spectral methods, and graph-based learning, where the structure of interactions among elements is essential for modeling global context [47, 48, 50].
Affinity matrices have long been used in machine learning and pattern recognition to represent graphs or networks of interactions. Early examples include graph adjacency matrices in spectral clustering (where $A_{ij}$ might be a Gaussian similarity between data points) [46, 47] and social network graphs for centrality analysis [48, 49].
In this work, we trace the origin of self-attention to a broader tradition of affinity-based reasoning. In particular, we highlight Infinite Feature Selection (Inf-FS), introduced in 2015, as a pivotal development that established the use of a fully connected affinity matrix $A$ to quantify and propagate feature importance [51]. Inf-FS provides a general-purpose framework where $A$ can be defined via statistical, semantic, or learned measures, and is used to:
  1. (1)
    Score features based on their connectedness;
  2. (2)
    Weight features according to their importance;
  3. (3)
    Rank features globally across a dataset;
  4. (4)
    Select informative subsets of features.
The key idea is that by comparing all pairs of elements, one can capture global structure or importance. The formulation is mathematically grounded in a power series expansion over $A$, which aggregates contributions from all possible paths of interaction between features—direct and indirect. This approach anticipated the key structural idea of self-attention: that relationships between input elements can be encoded in a pairwise affinity matrix and used to control the flow of information [51].
While Inf-FS was developed in the context of feature selection, the same structure, an affinity matrix $A$ used for weighted aggregation, appeared later in self-attention mechanisms, particularly in models where input elements (e.g., tokens, patches, nodes) are reweighted based on their contextual relevance. In this view, self-attention emerges as a special case where $A$ is constructed dynamically from learned input representations and applied in a single-hop update [61].
Self-attention (also called intra-attention) refers to an attention mechanism that relates a sequence or set with itself, as opposed to the earlier “attention” in sequence-to-sequence models which related one sequence with another (e.g., decoder attending to encoder states in machine translation [60]). Importantly, self-attention is implemented via an affinity matrix $A$ of size $N\times N$ (for $N$ input elements) that effectively forms a fully-connected graph over the elements. Each $A_{ij}$ indicates how much element $i$ should pay attention to element $j$. This structural idea—computing pairwise comparisons to derive contextual weights—originated in the domain of feature selection, particularly within the re-weighting step that precedes the selection of informative features [51, 58, 59]. In this setting, re-weighting serves to estimate the relative importance of each feature in relation to all others, based on affinity-based interactions. Although this process was not originally framed as “attention,” it anticipates the same mechanism of importance-driven weighting found in self-attention.
@@
This paper does not claim that Inf-FS introduced self-attention in its modern form. Rather, it emphasizes that Inf-FS introduced a general mathematical structure that was later adopted and operationalized in neural self-attention. Several later works, including AFS [34] and Sequential Attention [35], explicitly use attention weights for feature selection, further bridging these ideas.
We now present a historical timeline highlighting key developments in affinity-based reasoning across domains, culminating in the adoption of self-attention as a central mechanism in deep learning.
@@
The following timeline outlines major developments in the use of pairwise affinity matrices across disciplines, with an emphasis on their role in feature selection, attention mechanisms, and related learning frameworks.
  • 1950s–1960s (Psychological Foundations): Early theories of human attention emerged in cognitive psychology and neuroscience, including the cocktail party effect [52], Broadbent’s filter model [53], and Sperling’s partial report paradigm [54]. These works motivated later computational models of attention.
  • 1980s–1990s (Precursors to Computational Attention): Neural architectures incorporating multiplicative interactions, such as sigma-pi units [55], higher-order neural networks [56], and fast weight controllers [57], laid the groundwork for key-value mechanisms in attention.
  • 1998–2005 (Affinity in Image Processing and Graph Ranking): Bilateral filtering [7] and non-local means [8] used Gaussian affinity functions to smooth images based on patch similarity. Concurrently, the PageRank algorithm [9] ranked nodes in web graphs via random walks, effectively diffusing relevance through a pairwise affinity matrix. These methods demonstrated the utility of global interactions encoded via pairwise similarities, despite being non-learnable.
  • 2014 (Learned Attention in Feature Selection): Wang et al. introduced the “Attentional Neural Network” [19], incorporating top-down cognitive modulation for feature weighting in noisy visual recognition tasks. This was an early example of attention-inspired reweighting applied to feature selection.
  • 2015 (Infinite Feature Selection): Roffo et al. proposed Infinite Feature Selection (Inf-FS) [51], introducing a fully connected affinity matrix $A$ between features to model pairwise relationships. A ranking score is computed via a power series expansion:
@@
    Inf-FS formalized feature scoring, weighting, ranking, and selection through global affinity propagation. It did not frame the approach as “attention,” but it anticipated the structural form later used in self-attention mechanisms.
    Bahdanau et al. introduced attention for sequence-to-sequence models in neural machine translation [60]. Alignment scores were computed between encoder and decoder states, normalized via softmax to reweight source representations.
  • 2016–2017 (Self-Attention in Recurrent Models): Attention was extended within sequences via self-attention modules. Cheng et al. [20], Parikh et al. [22], and Lin et al. [23] applied intra-sequence attention for reading comprehension and sentence encoding, replacing or augmenting recurrent connections.
  • 2017 (Transformer Architecture): Vaswani et al. introduced the Transformer [61], which eliminated recurrence and adopted scaled dot-product self-attention as the core mechanism. Each layer computes:
@@
    This generalized the idea of contextual weighting via affinity matrices to end-to-end learning over sequences.
  • 2017–2019 (Pairwise Reasoning in Sets and Graphs): Relation Networks [24] modeled interactions over object pairs for relational reasoning. Set Transformers [25] used permutation-invariant self-attention for unordered sets. Graph Attention Networks (GAT) [31] introduced learnable attention over graph edges.
  • 2018 (Non-local Attention in Vision): Wang et al. proposed Non-Local Neural Networks [30], applying attention to image and video features by computing long-range dependencies via learned similarity functions, paralleling the Transformer formulation.
  • 2019–2020 (Efficient Attention): Several methods addressed the scalability of attention for long sequences. These include Reformer [32], Linformer [33], Performer [36], and Efficient Attention [37]. These variants introduced structural or kernel-based approximations to reduce computational cost while preserving pairwise weighting.
  • 2021 (Inf-FS and Structural Convergence): Ramsauer et al. [39] formalized connections between Hopfield networks and Transformers, interpreting attention as associative retrieval over affinity matrices.
  • 2022–2023 (Unified Affinity-Based Modeling): Self-attention mechanisms have since been adopted in vision (e.g., Vision Transformers), biology (e.g., AlphaFold), and vision-language models (e.g., CLIP). At the same time, several works applied attention weights to feature selection tasks explicitly, such as AFS [34] and Sequential Attention [35], reinforcing the continuity with the Inf-FS framework.
Although Inf-FS and self-attention were developed in different contexts—feature selection and sequence modeling, respectively—both frameworks employ a common structural element: a pairwise affinity matrix $A\in\mathbb{R}^{N\times N}$ that encodes relationships among input elements.
In Inf-FS, the matrix $A$ is used to model affinities between features and may be defined analytically (e.g., using correlation or distance metrics) or learned from data. The method ranks features by aggregating relevance scores over multi-hop paths on the induced affinity graph, where paths of arbitrary length are considered via a matrix power series. In contrast, self-attention computes $A$ dynamically from learned token embeddings, typically using a scaled dot-product formulation, and applies the resulting weights in a single-hop aggregation. Deeper interactions in self-attention arise through layer stacking rather than explicit multi-hop propagation.
Methodologically, Inf-FS provides a general framework for scoring and selecting elements based on affinity-based propagation. Self-attention aligns with this framework when the affinity matrix is constructed from data-dependent similarities and used to weight contextual contributions. The distinction lies in the construction and application of $A$: Inf-FS emphasizes global feature relevance across the input space, while self-attention focuses on local context within a specific forward pass.
This comparison highlights the structural similarity between the two approaches and illustrates how a common mathematical mechanism—pairwise affinity modeling—can be instantiated under different inductive biases and computational constraints.
@@ 2. Graph-Based Feature Selection and Early Affinity Approaches
[…]
@@
Feature Selection as Hard Attention: A Conceptual and Structural Justification
Feature selection and attention mechanisms both aim to allocate importance weights across inputs, yet they have historically emerged in different contexts, feature selection within classical machine learning and attention within neural architectures. Despite this, their underlying logic is closely aligned: identifying and emphasizing relevant inputs while suppressing less informative ones.
Traditionally, feature selection consists of three stages: (1) relevance estimation, where features are scored; (2) feature re-weighting, where feature magnitudes are modulated; and (3) selection, often by thresholding scores or applying sparsity constraints. Notably, it is the re-weighting stage—where feature importance is explicitly modeled via real-valued scores—that closely anticipates the logic of attention. In this sense, attention mechanisms can be interpreted as differentiable, context-aware feature re-weighting modules that generalize classical selection approaches.
This connection is further exemplified in modern architectures through the use of Feature Selection Gates (FSG) or Hard-Attention Gates (HAG), as introduced in [15, 16]. These gating mechanisms compute per-feature importance weights using learnable parameters, typically passed through a sigmoid activation. The resulting scores modulate input features via elementwise multiplication, structurally mirroring attention mechanisms that compute weighted sums of contextual representations. When paired with gradient routing (GR) [15], gating allows the separation of the feature selection process from downstream supervision, enabling independent optimization and interpretability of the selected subsets.
More broadly, attention mechanisms can be interpreted as a subclass of feature selection strategies—those that are:
  • Differentiable and trained end-to-end via backpropagation,
  • Instance-specific, adapting to each input sample,
  • Context-dependent, adjusting importance based on interactions among input components.
Traditional methods, such as Infinite Feature Selection (Inf-FS) [1], compute global importance scores prior to training. Inf-FS relies on an affinity matrix $A\in\mathbb{R}^{N\times N}$, capturing pairwise relationships among features, and derives importance scores via a matrix power series. While static and unsupervised, this process structurally resembles the affinity computation performed in attention mechanisms. Extensions of Inf-FS have further explored eigenvector centrality as a scoring metric [11], online feature selection for tracking [12], and biomedical applications such as stroke biomarker identification from neuroimaging data [13], demonstrating the generality of affinity-based feature reasoning.
Recent works have applied these ideas to deep models as well. In large language model pipelines, affinity-based reasoning has been leveraged to construct input-dependent feature filters and gates that reflect learned attention patterns [14]. These architectures explicitly integrate feature selection as part of their backbone, confirming that selection and attention are not isolated paradigms but mutually reinforcing. In practice, differentiable gating units also serve as drop-in replacements for static selection modules, offering both interpretability and adaptability.
Public implementations of classical feature selection methods like Inf-FS further facilitate this integration. The Feature Selection Library (FSLib) [17, 18] provides MATLAB-based tools for reproducible evaluation, while the Python package PyIFS supports integration into modern deep learning pipelines.
In this light, attention mechanisms do not replace traditional feature selection but rather extend them into a dynamic, supervised, and differentiable regime. Gates like FSG and mechanisms such as GR bridge these perspectives by preserving the core principle of importance-based weighting while adapting it to deep, task-specific architectures.
This framing also supports interpreting hard-attention gates as learnable and instance-specific approximations of classical subset selection—particularly when binarization or sparsity regularization is applied. Whereas soft attention distributes weight across all inputs, hard gates enforce a sparse support, closely aligned with the goal of selecting a meaningful feature subset.
In summary, the introduction of feature selection gates, gradient-aware gating strategies, and attention-based scoring in neural networks reflects a conceptual and structural continuity with prior work in pattern recognition. These developments reintroduce the foundational concepts of relevance, selection, and sparsity into scalable deep models—providing a common ground between classical feature selection and modern attention mechanisms.